OpenStudy (anonymous):

Wraite equation of the line through the given point with the given slope. Write an equation in slope-intercept form. (2,3);m=-2

OpenStudy (aroub):

y=mx+b , m=-2 , (2,3) 3=-2(2)+b 3=-4+b 7=b y=-2x+7

OpenStudy (amistre64):

y = mx -mPx + Py

OpenStudy (amistre64):

y = -2x +2(2)+3 y = -2x +4+3 y = -2x +7